# New mixed inhomogeneous phase in vortical gluon plasma: first-principle results from rotating SU(3) lattice gauge theory

### v2

@@ Abstract
Introduction.
The experimental observation of quark-gluon plasma (QGP) produced in non-central heavy-ion collisions has ignited intense theoretical interest in these highly vortical, strongly interacting systems. The plasma droplets are created in a state possessing a nonzero net angular momentum with the vorticity of the plasma estimated to reach the enormous value of Adamczyk et al. (2017)
@@ Introduction.
The rapid rotation, which is revealed experimentally via the global polarization measurements of produced hyperons Abelev et al. (2007); Adamczyk et al. (2017); Becattini and Lisa (2020), is analytically predicted to have a significant impact on the phase structure of QGP Chen et al. (2016); Chernodub and Gongyo (2017); Wang et al. (2019); Zhang et al. (2020); Sadooghi et al. (2021); Fujimoto et al. (2021); Chen et al. (2022); Zhao et al. (2023); Chernodub (2021); Mameda and Takizawa (2023); Sun et al. (2023); Satapathy (2023); Braga and Junqueira (2024); Eto et al. (2023a, b); Jiang (2023).
@@
It must be noted that all theoretical approaches to the QGP thermodynamics are performed under the condition of the rigid rotation of the system Ambruş and Winstanley (2014); Chen et al. (2016); Chernodub and Gongyo (2017); Wang et al. (2019); Zhang et al. (2020); Sadooghi et al. (2021); Fujimoto et al. (2021); Golubtsova et al. (2022); Chen et al. (2022); Zhao et al. (2023); Chernodub (2021); Mameda and Takizawa (2023); Sun et al. (2023); Satapathy (2023); Braga and Junqueira (2024); Eto et al. (2023a, b); Jiang (2023); Braguta et al. (2020, 2021, 2022, 2023a, 2023b); Yang and Huang (2023). In a rigidly rotating fluid, the uniformity of vorticity implies the absence of the viscous component of the stress tensor Landau and Lifshitz (1982), indicating no internal friction and no entropy production, thereby representing a thermodynamic ground state of the system. Although this simplification is not an accurate approximation of an off-equilibrium state of the QGP produced in heavy-ion collisions, it drastically simplifies the theoretical analysis and makes predictions on the consequences of rotating QCD much easier. Despite the would-be simplicity of the rigid rotation, it reveals profound disagreement between the results coming from analytical Chen et al. (2016); Chernodub and Gongyo (2017); Wang et al. (2019); Zhang et al. (2020); Sadooghi et al. (2021); Fujimoto et al. (2021); Chen et al. (2022); Zhao et al. (2023); Chernodub (2021); Mameda and Takizawa (2023); Sun et al. (2023); Satapathy (2023); Braga and Junqueira (2024) and numerical Braguta et al. (2020, 2021, 2022, 2023a, 2023b); Yang and Huang (2023) simulations (see also Jiang (2023)), implying that our understanding of the rotating QGP is far from satisfactory.
@@
The controversy is further augmented by two independent numerical observations in finite-temperature Yang-Mills theory, made in both rotating Braguta et al. (2023b) and static Braguta et al. (2023c) systems, which revealed that the moment of inertia of the gluon plasma takes a negative value in a range of temperatures starting from a temperature slightly below the deconfining phase transition $T_{c0}\equiv T_{c}(\Omega=0)$, identified in a non-rotating system, up to the “supervortical” temperature $T_{s}\simeq 1.5T_{c0}$ Braguta et al. (2023b, c, d). This exceptional phenomenon, which has no analogs in ordinary fluids, was shown to occur due to thermal evaporation of the chromomagnetic condensate Braguta et al. (2023b). The negative moment of inertia was argued to emerge from the negativity of a gluonic Barnett effect, which implies a negative coupling of the gluon spin polarization with vorticity Braguta et al. (2023e).
@@
Below, we deepen this mystery even further. We reveal, using first-principle lattice simulations, a new mixed structure of confining (hadron) and deconfining (quark-gluon plasma) phases that co-exist simultaneously in a rotating hot gluon matter. Although the existence of a confining-deconfining phase was proposed some years ago Chernodub (2021) (see also the holographic study Braga and Junqueira (2024)) and was evidenced in lattice calculations with a kinematically defined order parameter Chernodub et al. (2022) we show that the dynamical structure of the plasma is surprisingly different from all existing theoretical predictions.
Lattice set up for rotating gluodynamics.
@@
Rigid rotation with a fixed angular velocity $\Omega$ can be conveniently described by setting a relevant system in curved spacetime endowed with the metric Yamamoto and Hirono (2013):
@@
To study this system, we follow lattice methods used in Refs. Yamamoto and Hirono (2013); Braguta et al. (2020, 2022, 2021, 2023a, 2023b, 2023c, 2023d, 2023e). In particular, within lattice simulations, one studies the system with the action (3) in thermodynamic equilibrium. The partition function of this system can be written as the path integral over gluon fields with the weight $e^{-S_{E}}$, where $S_{E}$ is the Euclidean lattice action of rotating Yang-Mills theory. Unfortunately, this action is a complex-valued function that leads to a sign problem and does not allow to apply lattice methods directly to this theory. To overcome this obstacle, we carry out lattice simulation with imaginary angular velocity $\Omega_{I}=-i\Omega$ and then analytically continue the results to real values of $\Omega$ through the identification: $\Omega_{I}^{2}\to-\Omega^{2}$.
The temperature of a system in the thermodynamic equilibrium in an external static gravitational field depends on the coordinates $\boldsymbol{r}$ and obeys the well-known Tolman-Ehrenfest (TE) law Tolman (1930); Tolman and Ehrenfest (1930): $\sqrt{g_{00}({\boldsymbol{r}})}T({\boldsymbol{r}})=T_{0}=\mathrm{const}$. For a rotating system (2), the TE law gives:
where $T_{0}$ is the temperature at the rotation axis ($r=0$). The last equality in Eq. (4) corresponds to the case of imaginary rotation. In our numerical simulations, we fix the on-axis temperature parameter $T_{0}$ and call it, unless stated otherwise, the temperature $T$ of gluon plasma.
Lattice setup.
@@
To construct the lattice form of the continuum action (3), we discretize the terms, which are coupled with the angular velocity, following Refs. Yamamoto and Hirono (2013); Braguta et al. (2021), whereas for the remaining (non-rotating) terms, we use tree-level improved Symanzik gauge action Curci et al. (1983); Luscher and Weisz (1985). The explicit form of the used lattice action may be found in Ref. Braguta et al. (2023b). The simulations were performed on the lattices of size $N_{t}\times N_{z}\times N_{s}^{2}$ ($N_{x}=N_{y}=N_{s}$), and the rotation axis passes through the lattice site in the center of $xy$-plane. To study in detail the non-trivial dynamics of gluons induced by the rotation, we use the lattices with a large extension in transversal space directions: $4\times 24\times 145^{2}$, $5\times 30\times 181^{2}$ and $6\times 36\times 217^{2}$.
@@
We applied open boundary conditions Braguta et al. (2021) in the directions that are orthogonal to the rotation axis and used conventional periodic boundary conditions in the directions along the rotation axis and the Euclidean time. While the choice of the boundary conditions in transversal directions warrants some discussion, the previous study Braguta et al. (2021) indicates that a type of spatial boundary does not produce an important impact on the bulk of the system due to strong screening. Nonetheless, to validate the correctness of this approach, we reproduced some of our results with periodic boundary conditions in $xy$-plane.
Slow rotation and analytic continuity.
@@
We carry out simulations at imaginary angular frequencies and then perform analytic continuation to real rotations. It is convenient to describe the results in terms of the imaginary velocity $v_{I}=\Omega_{I}R$ at the QGP boundary, $v_{I}^{2}=-v_{R}^{2}$, where the velocity $v_{I}$ is taken for the middle of the boundary side, i.e., at the distance $R=a(N_{s}-1)/2$ from the axis of rotation. The analytic continuation is only possible if at any point of the system $|\Omega_{I}|r<1$ so that the real solid rotation does not violate causality with a corresponding theorem proven in Ref. Ambruş and Chernodub (2023). In our paper, we respect the causality bound.
@@
Another important and, perhaps, not intuitively clear statement is that the thermodynamical properties of the rotating plasma, which has the shape of a cylinder, depend on the angular frequency $\Omega$ only in the combination $v_{R}=\Omega R$ which is nothing but the (imaginary) velocity of the system at the boundary $v_{R}$ Braguta et al. (2023b, c, e).
Confining and deconfining phases.
@@
Figure 1: (top) The distribution of the local Polyakov loop in $x,y$-plane for lattice of size $5\times 30\times 181^{2}$ with open boundary conditions at the fixed on-axis temperature $T=0.95\,T_{c0}$ and different imaginary angular frequencies (also shown as imaginary velocities at the boundary, $v_{I}^{2}\equiv(\Omega_{I}R)^{2}=0.04,0.12,0.24,0.48$) with $R=13.5$ fm. (bottom) The Polyakov loop at the $x$ axis. The vertical lines mark the phase boundaries with shaded uncertainties. The violet (blue) data points correspond to periodic (open) boundary conditions. Movies on the phase evolution with increasing $\Omega_{I}$ are available as ancillary files Braguta et al. (2023f).
@@
Finalizing this section, we stress that the central confining regions in Fig. 1 have the form of a disk, despite the lattice having a square shape, thus signaling the expected restoration of the rotational symmetry and implying that we work in the physical domain of lattice coupling close to the continuum limit. Moreover, the boundary conditions affect the local phase structure only very near the boundary. The latter property is a result of the short-range nature of the screening, which implies that the boundary effects on the phase structure are negligible Braguta et al. (2021).
Size of the inhomogeneity.
@@
At a fixed distance $r$ from the rotation axis, the expectation value of the Polyakov loop and its susceptibility can only be evaluated at a finite number of spatial points proportional to the lattice extension $L_{z}$. Since our calculations are performed at finite $L_{z}$, a small volume of this lattice submanifold leads to high uncertainty in the determination of the critical temperature. To reduce the associated statistical error, we calculated the mentioned quantities within a thin cylinder $(r-\delta r/2,r+\delta r/2)$. We justified our approach by demonstrating numerically that the finiteness of $\delta r$ brings only a minor systematic error to the estimation of the critical temperature Braguta et al. (2024).
@@
where the transition temperature on the rotation axis, $T_{c}$, and the dimensionless “vortical curvature” $\kappa$  [^2] The vortical curvature $\kappa$ resembles the finite-density curvature of the QCD phase transition at small values of the baryonic chemical potential de Forcrand and Philipsen (2002). serve as the fitting parameters. The best fits for various angular frequencies are shown in the main plot of Fig. 2.
On-axis transition and vortical curvature.
@@
Our result is even more surprising given that the critical temperature of the deconfining transition in all analytic calculations available so far is predicted to exhibit a significant dependence on rotation Chen et al. (2016); Chernodub and Gongyo (2017); Wang et al. (2019); Zhang et al. (2020); Sadooghi et al. (2021); Fujimoto et al. (2021); Chen et al. (2022); Zhao et al. (2023); Chernodub (2021); Mameda and Takizawa (2023); Sun et al. (2023); Satapathy (2023); Braga and Junqueira (2024); Jiang (2023). In addition, the previous numerical results that have found a dependence of the critical temperature on $\Omega_{I}$ without specifying the distance of the center of rotation Braguta et al. (2020, 2021, 2022, 2023a, 2023b); Yang and Huang (2023) should be understood as the bulk-averaged results. We found a minor dependence of our results on the lattice spacing and obtained a value $\kappa=0.901(32)$ after continuum limit extrapolation using the data for $N_{t}=4,5,6$.
Violation of the Tolman-Ehrenfest law.
The TE law (4) suggests that real rotation effectively heats the system outside of the rotation axis. This fact led Ref. Chernodub (2021) to conclude –also supported by a calculation in a low-dimensional confining model– that the outer region of the rotating plasma is hotter than its interior (thus, with $\left\langle L\right\rangle\neq 0$ far from the axis and $\left\langle L\right\rangle=0$ close to the axis), which naturally leads to the two-phase structure of the rotating plasma in thermal equilibrium. For imaginary rotation, the TE law is represented by the second equality of the equation (4), which leads to the following two-phase structure: deconfinement at the center and confinement close to the boundaries. This theoretically motivated phase structure is in contradiction with the numerical simulations of this paper. Thus the TE law is violated in rotating gluodynamics.
The Tolman-Ehrenfest law is violated since external gravity, generated by centrifugal forces, influences the dynamics of gluon matter and cannot be accounted for by the simple formula (4). We suggest that this effect appears due to conformal anomaly, which is known to affect the temperature distribution in background gravitational field Gim and Kim (2015); Eune and Kim (2017). This suggestion aligns well with the indication that a condensed-matter counterpart of the TE law, the Luttinger relation Luttinger (1964), is violated in certain materials Park et al. (2022); Bermond et al. (2022) as a result of the conformal anomaly Bermond et al. (2022). Strikingly, it is the conformal anomaly, which, according to Refs. Braguta et al. (2023c, b), might be responsible for the negative moment of inertia of the gluon plasma Braguta et al. (2023e) so that these exotic effects can be related to each other.
@@
Phase diagram for imaginary and real rotation.
We construct the phase diagram in the $(\Omega,r)$ plane using the results for local critical temperature extrapolated to the continuum limit. The radius of the spatial boundary between the confining phase close to the axis and the deconfining phase in an outer region depends on temperature and angular frequency, as shown in Fig. 3. As the rotation increases, the radius separating these phases diminishes. This tendency is observed at all studied temperatures for imaginary rotation. The radius of the phase boundary shrinks to zero when the temperature at the center of the rotating plasma reaches the deconfining temperature of the non-rotating plasma.
For the studied range of temperatures and angular frequencies, the phase diagram, shown in Fig. 3, is precisely the same for imaginary and real rotations up to the swap of the phases. This property follows from the quadratic form of the critical curve (6), with the swap of the phases explained in detail in Fig. 4.
The role of fermions.
While dynamical fermions affect the thermodynamics of QGP, it is the dynamics of gluons that makes the plasma a perfect fluid possessing an exceptionally low viscosity Bernhard et al. (2019) and determines the hydrodynamics of its evolution. Thus, we expect that the observed inhomogeneous features will also persist in QGP.
Summary and Conclusions.

### v3

@@ Abstract
keywords:
Relativistic rotation , Quark-gluon plasma , Phase transition , Lattice QCD , Heavy-ion collisions , Mixed phase
journal: Physics Letters B
[…]
1 Introduction
The experimental observation of quark-gluon plasma (QGP) produced in non-central heavy-ion collisions has ignited intense theoretical interest in these highly vortical, strongly interacting systems. The plasma droplets are created in a state possessing a nonzero net angular momentum with the vorticity of the plasma estimated to reach the enormous value of [1]
@@ 1 Introduction
The rapid rotation, which is revealed experimentally via the global polarization measurements of produced hyperons [2, 1, 3], is analytically predicted to have a significant impact on the phase structure of QGP [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19].
@@
It must be noted that all theoretical approaches to the QGP thermodynamics are performed under the condition of the rigid rotation of the system [20, 4, 5, 6, 7, 8, 9, 21, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 22, 23, 24, 25, 26, 27]. In a rigidly rotating fluid, the uniformity of vorticity implies the absence of the viscous component of the stress tensor [28], indicating no internal friction and no entropy production, thereby representing a thermodynamic ground state of the system. Although this simplification is not an accurate approximation of an off-equilibrium state of the QGP produced in heavy-ion collisions, it drastically simplifies the theoretical analysis and makes predictions on the consequences of rotating QCD much easier. Despite the would-be simplicity of the rigid rotation, it reveals profound disagreement between the results coming from analytical [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16] and numerical [22, 23, 24, 25, 26, 27] simulations (see also [19]), implying that our understanding of the rotating QGP is far from satisfactory.
@@
The controversy is further augmented by two independent numerical observations in finite-temperature Yang-Mills theory, made in both rotating [26] and static [29] systems, which revealed that the moment of inertia of the gluon plasma takes a negative value in a range of temperatures starting from a temperature slightly below the deconfining phase transition $T_{c0}\equiv T_{c}(\Omega=0)$, identified in a non-rotating system, up to the “supervortical” temperature $T_{s}\simeq 1.5T_{c0}$ [26, 29, 30]. This exceptional phenomenon, which has no analogs in ordinary fluids, was shown to occur due to thermal evaporation of the chromomagnetic condensate [26]. The negative moment of inertia was argued to emerge from the negativity of a gluonic Barnett effect, which implies a negative coupling of the gluon spin polarization with vorticity [31].
@@
Below, we deepen this mystery even further. We reveal, using first-principle lattice simulations, a new mixed structure of confining (hadron) and deconfining (quark-gluon plasma) phases that co-exist simultaneously in a rotating hot gluon matter. Although the existence of a confining-deconfining phase was proposed some years ago [12] (see also the holographic study [16]) and was evidenced in lattice calculations with a kinematically defined order parameter [32] we show that the dynamical structure of the plasma is surprisingly different from all existing theoretical predictions.
2 Rotating gluodynamics on the lattice
@@
Rigid rotation with a fixed angular velocity $\Omega$ can be conveniently described by setting a relevant system in curved spacetime endowed with the metric [33]:
@@
To study this system, we follow lattice methods used in Refs. [33, 22, 24, 23, 25, 26, 29, 30, 31]. In particular, within lattice simulations, one studies the system with the action (3) in thermodynamic equilibrium. The partition function of this system can be written as the path integral over gluon fields with the weight $e^{-S_{E}}$, where $S_{E}$ is the Euclidean lattice action of rotating Yang-Mills theory. Unfortunately, this action is a complex-valued function that leads to a sign problem and does not allow to apply lattice methods directly to this theory. To overcome this obstacle, we carry out lattice simulation with imaginary angular velocity $\Omega_{I}=-i\Omega$ and then analytically continue the results to real values of $\Omega$ through the identification: $\Omega_{I}^{2}\to-\Omega^{2}$.
@@
To construct the lattice form of the continuum action (3), we discretize the terms, which are coupled with the angular velocity, following Refs. [33, 23], whereas for the remaining (non-rotating) terms, we use tree-level improved Symanzik gauge action [34, 35]. The explicit form of the used lattice action may be found in Ref. [26]. The simulations were performed on the lattices of size $N_{t}\times N_{z}\times N_{s}^{2}$ ($N_{x}=N_{y}=N_{s}$), and the rotation axis passes through the lattice site in the center of $xy$-plane. To study in detail the non-trivial dynamics of gluons induced by the rotation, we use the lattices with a large extension in transversal space directions: $4\times 24\times 145^{2}$, $5\times 30\times 181^{2}$ and $6\times 36\times 217^{2}$.
In imaginary time formalism, the temperature of the system equals the inverse period along the temporal direction on the lattice. In order to avoid the particularities in the definition of temperature in the gravitational background (see Sec. 8), we denote by $T$ the value of temperature defined precisely on the rotation axis [22, 23], where the metric-induced effects vanish, unless stated otherwise.
@@
We applied open boundary conditions [23] in the directions that are orthogonal to the rotation axis and used conventional periodic boundary conditions in the directions along the rotation axis and the Euclidean time. While the choice of the boundary conditions in transversal directions warrants some discussion, the previous study [23] indicates that a type of spatial boundary does not produce an important impact on the bulk of the system due to strong screening. Nonetheless, to validate the correctness of this approach, we reproduced some of our results with periodic boundary conditions in $xy$-plane.
3 Slow rotation and analytic continuity
@@
We carry out simulations at imaginary angular frequencies and then perform analytic continuation to real rotations. It is convenient to describe the results in terms of the imaginary velocity $v_{I}=\Omega_{I}R$ at the QGP boundary, $v_{I}^{2}=-v_{R}^{2}$, where the velocity $v_{I}$ is taken for the middle of the boundary side, i.e., at the distance $R=a(N_{s}-1)/2$ from the axis of rotation. The analytic continuation is only possible if at any point of the system $|\Omega_{I}|r<1$ so that the real solid rotation does not violate causality with a corresponding theorem proven in Ref. [36]. In our paper, we respect the causality bound.
@@
Another important and, perhaps, not intuitively clear statement is that the thermodynamical properties of the rotating plasma, which has the shape of a cylinder, depend on the angular frequency $\Omega$ only in the combination $v_{R}=\Omega R$ which is nothing but the (imaginary) velocity of the system at the boundary $v_{R}$ [26, 29, 31]. In the case of slow rotation, this statement follows directly from the extensive properties of the moment of inertia, and it was numerically confirmed in lattice gluodynamics in Refs. [29, 26] using two independent methods of calculations. For a rotating system of free bosons, it was also proven analytically in Ref. [36].
4 Confining and deconfining phases
@@
Figure 1: (top) The distribution of the local Polyakov loop in $x,y$-plane for lattice of size $5\times 30\times 181^{2}$ with open boundary conditions at the fixed on-axis temperature $T=0.95\,T_{c0}$ and different imaginary angular frequencies (also shown as imaginary velocities at the boundary, $v_{I}^{2}\equiv(\Omega_{I}R)^{2}=0.04,0.12,0.24,0.48$) with $R=13.5$ fm. (bottom) The Polyakov loop at the $x$ axis. The vertical lines mark the phase boundaries with shaded uncertainties. The violet (blue) data points correspond to periodic (open) boundary conditions. Movies on the phase evolution with increasing $\Omega_{I}$ are available as ancillary files [37].
@@ 5 Emergence of the inhomogeneity
Finalizing this section, we stress that the central confining regions in Fig. 1 have the form of a disk, despite the lattice having a square shape, thus signaling the expected restoration of the rotational symmetry and implying that we work in the physical domain of lattice coupling close to the continuum limit. Moreover, the boundary conditions affect the local phase structure only very near the boundary. The latter property is a result of the short-range nature of the screening, which implies that the boundary effects on the phase structure are negligible [23].
6 Size of the inhomogeneity
@@
At a fixed distance $r$ from the rotation axis, the expectation value of the Polyakov loop and its susceptibility can only be evaluated at a finite number of spatial points proportional to the lattice extension $L_{z}$. Since our calculations are performed at finite $L_{z}$, a small volume of this lattice submanifold leads to high uncertainty in the determination of the critical temperature. To reduce the associated statistical error, we calculated the mentioned quantities within a thin cylinder $(r-\delta r/2,r+\delta r/2)$. We justified our approach by demonstrating numerically that the finiteness of $\delta r$ brings only a minor systematic error to the estimation of the critical temperature [38].
@@
where the transition temperature on the rotation axis, $T_{c}$, and the dimensionless “vortical curvature” $\kappa$ serve as the fitting parameters. [^2] The vortical curvature $\kappa$ resembles the finite-density curvature of the QCD phase transition at small values of the baryonic chemical potential [39]. The best fits for various angular frequencies are shown in the main plot of Fig. 2.
7 On-axis transition and vortical curvature
@@
Our result is even more surprising given that the critical temperature of the deconfining transition in all analytic calculations available so far is predicted to exhibit a significant dependence on rotation [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 19]. In addition, the previous numerical results that have found a dependence of the critical temperature on $\Omega_{I}$ without specifying the distance of the center of rotation [22, 23, 24, 25, 26, 27] should be understood as the bulk-averaged results. We found a minor dependence of our results on the lattice spacing and obtained a value $\kappa=0.902(33)$ after continuum limit extrapolation using the data for $N_{t}=4,5,6$.
8 Violation of the Tolman-Ehrenfest law
The temperature of a system in the thermodynamic equilibrium in an external static gravitational field depends on the coordinates $\boldsymbol{r}$ and obeys the well-known Tolman-Ehrenfest (TE) law [40, 41]: $\sqrt{g_{00}({\boldsymbol{r}})}T({\boldsymbol{r}})=T_{0}=\mathrm{const}$. For a rotating system (2), the TE law gives:
where $T_{0}$ is the temperature at the rotation axis ($r=0$). The last relation in Eq. (6) corresponds to the case of imaginary rotation. To simplify notations, we use the on-axis temperature $T_{0}\equiv T$, Eq. (6), to refer to the temperature of the gluon plasma.
The TE law (6) suggests that real rotation effectively heats the system outside of the rotation axis. This fact led Ref. [12] to conclude—also supported by a calculation in a low-dimensional confining model—that the outer region of the rotating plasma is hotter than its interior (thus, with $\left\langle L\right\rangle\neq 0$ far from the axis and $\left\langle L\right\rangle=0$ close to the axis), which naturally leads to the two-phase structure of the rotating plasma in thermal equilibrium. For imaginary rotation, the TE law is represented by the second equality of the equation (6), which leads to the following two-phase structure: deconfinement at the center and confinement close to the boundaries. This theoretically motivated phase structure is in contradiction with the numerical simulations of this paper. Thus, the TE law is violated in rotating gluodynamics.
The Tolman-Ehrenfest law is violated since external gravity, generated by centrifugal forces, influences the dynamics of gluon matter and cannot be accounted for by the simple formula (6). We suggest that this effect appears due to conformal anomaly, which is known to affect the temperature distribution in background gravitational field [42, 43]. This suggestion aligns well with the indication that a condensed-matter counterpart of the TE law, the Luttinger relation [44], is violated in certain materials [45, 46] as a result of the conformal anomaly [46]. Strikingly, it is the conformal anomaly, which, according to Refs. [29, 26], might be responsible for the negative moment of inertia of the gluon plasma [31] so that these exotic effects can be related to each other.
@@ 8 Violation of the Tolman-Ehrenfest law
9 Phase diagram for imaginary and real rotation
The mixed confinement-deconfinement phase may appear or not, depending on the specific values of the temperature $T$, angular velocity $\Omega$ and system radius $R$. In order to visualize the conditions for the co-existence of phases, we construct the phase diagram in the $(\Omega,r)$ plane using the results for local critical temperature extrapolated to the continuum limit.333 In constructing the diagram, Fig. 3, we assumed that $T_{c}(0)/T_{c0}=1$ at $v=0$ and included minor difference between $T_{c}(0)/T_{c0}$ in our lattice data and unity into uncertainty. The radius $r$ of the spatial boundary between the confining phase close to the axis and the deconfining phase in an outer region depends on temperature and angular frequency, as shown in Fig. 3. As the rotation increases, the radius separating these phases diminishes. This tendency is observed at all studied temperatures for imaginary rotation. The radius of the phase boundary shrinks to zero when the temperature at the center of the rotating plasma reaches the deconfining temperature of the non-rotating plasma.
For the studied range of temperatures and angular frequencies, the phase diagram, shown in Fig. 3, is precisely the same for imaginary and real rotations up to the swap of the phases. In Fig. 4 we graphically describe how the phase diagram shown in Fig. 3 is built. The green straight line in Fig. 4 represents the dependence of the local critical temperature on the local velocity of rotational motion $\Omega r$, given by Eq. (5). The region of imaginary rotation, $(\Omega r)^{2}<0$, is available for lattice simulation, whereas the results for real angular velocity, $(\Omega r)^{2}>0$, are obtained after the analytic continuation. For a given value of the temperature $T=T_{c0}-\Delta T$ the shape of the boundary between two phases on the $(\Omega_{I},r)$-plane (left inset in Fig. 4) follows from the condition: $T_{c}(\Omega_{I}r)=T_{c0}-\Delta T$. Graphically, this condition is represented by the intersection of the dashed line $T=T_{c0}-\Delta T=const$ with the solid line $T_{c}(\Omega_{I}r)$. For a symmetric value of the temperature $T=T_{c0}+\Delta T$ the critical condition is quite similar, $T_{c}(\Omega r)=T_{c0}+\Delta T$, but the angular velocity is real now (the resulting diagram is shown in the right inset in Fig. 4). The quadratic form of the critical curve (5) ensures that the phase diagrams for these two symmetric values of the temperature $T=T_{c0}\pm\Delta T$ have the same form up to the swap of the phases and the substitution of the imaginary angular velocity to the real one. Note that this identity of the phase diagrams for real and imaginary rotation may be violated by high order $\mathcal{O}(\Omega^{4})$-terms in Eq. (5), which are beyond the scope of this study.
10 The role of fermions
In the paper [25], the influence of relativistic rotation on the QCD critical temperatures averaged over the lattice volume was studied. In particular, the authors addressed the question of how rotation influences the gluon and quark sectors of QCD. They found that rotation applied only to the gluon sector of QCD enhances the critical temperatures. On the contrary, rotation of the quark sector only decreases the critical temperatures. If one rotates both sectors simultaneously, the critical temperature increases. In addition, it was found that the quark sector of QCD plays a minor role in changing the critical temperatures due to the rotation. For this reason, we expect that the observed inhomogeneous features will also persist in QGP.
11 Summary and Conclusions